Cut fruit is packed for sale in a supermarket. The weights of packages are normally distributed with a mean of 201.49 grams and variance of 24.63 grams.

Given that 5% of the distribution lies below the Z value of -1.64485, find the package weight that has 5% of the distribution above it. Give your result correct to four decimal places.


1
Expert's answer
2021-02-24T06:58:15-0500

Given "P(Z<-1.64485)=0.05." Then "P(Z>1.64485)=0.05."


"Z=\dfrac{X-\mu}{\sigma}"

Given "\mu=201.49\ g, \sigma=24.63\ g."

Then


"\dfrac{X-201.49}{24.63}=1.64489"

"X=1.64489(24.63)+201.49\approx242.0036"

"242.0036" grams.